Three branches, possessing a resistance of 50ohm, an inductance of 0.15H and a capacitor of 100µF respectively, are connected in parallel across a 100V, 50Hz supply. Calculate


a) The current in each branch


b) The supply current


  • c) The phase angle between the supply current and the supply voltage.
1
Expert's answer
2021-11-25T16:23:16-0500

a)

"I_R=U\/R=100\/50=2" A


"I_L=\frac{U}{2\pi fL}=\frac{100}{2\pi \cdot50\cdot 0.15}=2.12" A


"I_C=2\pi fCU=2\pi \cdot50\cdot100\cdot10^{-6}\cdot100=3.14" A


b)

"I=\sqrt{I_R^2+(I_C-I_L)^2}=\sqrt{2^2+(3.14-2.12)^2}=2.24" A


c)

"cos\phi=I_R/I=2/2.24=0.893"

"\phi=26\degree"

Since IC​ is greater than IL​ , the supply current leads the supply voltage by "26\degree"